LadyRavyRay:

Give the slope of the graph and the unit rate. Express your answers as simplified fractions, if necessary. Jorge: 2 miles every 3 hours The slope of the graph is His unit rate is

LadyRavyRay:

snowflake0531:

Slope is rise over run, so find another coordinate that isn't the origin

snowflake0531:

e.e let's just take (2,3) then, iwht the origin, so we see that it goes up 2, and to the right 3, and slope is rise over run so the slope is .... __ Then for unit rate, it's the same thing as slope in this case

LadyRavyRay:

Um..I'm still VERY confused

LadyRavyRay:

@snowflake0531

dude:

Pick a point on the graph

dude:

Although (2,3) is not a point on this graph, you can still use another point In this case I used (6,4) Going back to what snowflake said, the slope is \(\dfrac{"rise"}{"run"}\)

1 attachment
dude:

Rise being how much it goes "up" or "down" (Up is positive and down is negative) Run being how much it goes "left" or "right" (Left is negative, right is positive) In this case it goes up 4 units up and 6 units to the right \(\dfrac{4}{6}\) (You can simplify this but does it make sense?)

LadyRavyRay:

yeah

LadyRavyRay:

Is the unit rate, 1.5?

LadyRavyRay:

@dude is the answer for the unit rate, 1.5?

dude:

The slope is the same as the rate It is not 1.5 s: (Remember: Use \(\dfrac{4}{6}\))